# Observability-Aware Control for Quadrotor Formation Flight with Range-only Measurement

### v2

@@ Appendix A Formulas regarding likelihood ratios of infinite-dimensional spaces
whenever $\mathcal{K}^{-1}h\in L^{2}([0,T])$. Comparing Eqs. (56) and (57) shows that, for any integer $s>\alpha+1$, the condition $h\in\norm{h}_{H^{s}([0,T])}$ is sufficient for the right-hand side of Eq. (57) to converge, and moreover in such a case an inequality of the form Eq. (54) holds.
Remark.
An analogous construction to Example A.1 also works for the Fourier basis instead of the Legendre basis on $L^{2}([0,T])$, but one has to restrict to periodic functions when considering the Sobolev and Fréchet spaces on $[0,T]$.
Appendix B Asymptotic size of the minimum eigenvalue of the STLOG
We derive Proposition 5.2 in this appendix. This asymptotic result relates the STLOG in Eq. (34) to the local observability index $r_{*}$ (14) defined as a bridge to rank-testing observability in Sec. 3. We consider a fixed operating point of the system $(\mathbf{x},\mathbf{u})$, and adopt several inessential simplifications. Specifically, we set $\boldsymbol{\Sigma}=\mathbf{1}$, and simply denote the order-$r$ STLOG from Eq. (34) by $\mathbf{W}^{(r)}$ without the subscript $\boldsymbol{\Sigma}^{-1}$.
@@ Appendix B Asymptotic size of the minimum eigenvalue of the STLOG
We prove Part 1 in Proposition B.3, and provide expressions for $C_{1}(\mathbf{x},\mathbf{u}),C_{2}(\mathbf{x},\mathbf{u})$ in Propositions B.4 and B.7 respectively. ∎
In Section 3, we reviewed that the system is locally weakly observable at order $r$ if $\text{rank }\boldsymbol{\mathcal{O}}^{(r)}=\dim\mathcal{X}$. We introduce the following subspace:
@@
Clearly, $\mathcal{H}_{k-1}\supset\mathcal{H}_{k}$. This decreasing sequence of subspaces reaches $\{0\}$ precisely at the local observability index $r_{*}$. We collect some elementary consequences of this definition into the following Lemma:
@@
We rewrite the STLOG from Eq. (34) in terms of $\boldsymbol{\mathcal{O}}^{(r)}$ before evaluating the integral (27). For any $\mathbf{v},\tilde{\mathbf{v}}\in\mathcal{X}$, [^13] In Eq. (60), if instead of the standard inner product on $\mathcal{Y}$ we use the inner product induced by $\boldsymbol{\Sigma}^{-1}$, we recover the expressions in Eqs. (27) and (34).
@@
making $\mathbf{W}^{(r)}$ positive semi-definite by construction. This is linked to local observability by Proposition B.3.
@@
Next, we derive the upper and lower bounds in Eq. (B.1) in Theorem B.1. Let $r_{*}$ be the local observability index. We only consider $\mathbf{W}^{(r)}$ with $r\geq r_{*}$ in the following, and assume $r_{*}>0$ without loss of generality.
We first derive the upper bound in Eq. (B.1). We leverage the identity on analytical minima of a quadratic form over a unit sphere $\lambda_{\min}(\mathbf{W}^{(r)})=\min_{\mathbf{v}\in\mathcal{X},\ \norm{\mathbf{v}}=1}\mathbf{v}^{\top}\mathbf{W}^{(r)}\mathbf{v}$. We thus seek an appropriate unit vector $\mathbf{v}$ that demonstrates the upper bound in Eq. (B.1). Any unit vector $\mathbf{v}\in\mathcal{H}_{r^{*}-1}$ satisfies
@@
$\norm{(DL_{\mathbf{f}}^{r_{*}}\mathbf{h})\mathbf{v}}=\sigma_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}|_{\mathcal{H}_{r^{*}-1}}\right)>0,$ (63)
where $|_{\mathcal{H}_{r^{*}-1}}$ denotes the restriction to the subspace $\mathcal{H}_{r^{*}-1}$. Now we evaluate $\mathbf{v}^{\top}\mathbf{W}^{(r)}\mathbf{v}$ using our selected $\mathbf{v}$:
@@
$\lambda_{\min}(\mathbf{W}^{(r)})\leq\frac{(1+\epsilon)\Delta T^{2r_{*}+1}\sigma^{2}_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}|_{\mathcal{H}_{r^{*}-1}}\right)}{{(r_{*}!)}^{2}(2r_{*}+1)},$ (67)
@@
Next we derive the lower bound in (58). It is convenient to rewrite the STLOG (34) into block matrix form
@@
$\alpha\sigma^{2}_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}\right)\leq\frac{\lambda_{\min}(\mathbf{W}^{(r)})}{\Delta T^{2r_{*}+1}}\leq\beta\sigma^{2}_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}|_{\mathcal{H}_{r^{*}-1}}\right),$ (74)
where $\alpha,\beta$ are constants. Thus, we proved the inequality (58) in Theorem B.1 by explicitly constructing $C_{1}(\mathbf{x},\mathbf{u}),C_{2}(\mathbf{x},\mathbf{u})$ in terms of $\boldsymbol{\mathcal{O}}^{(r_{*})}$ only. The last part of Theorem B.1 follows from the Bolzano—Weierstrass theorem.
Appendix C Choice of STLOG order for quadrotor CLS
@@
$\begin{bmatrix}\dot{\mathbf{z}}\\ \dot{\mathbf{q}}\end{bmatrix}$ $=\begin{bmatrix}\mathbf{f}_{z}(\mathbf{z},\mathbf{q},\mathbf{u})\\ \mathbf{f}_{q}(\mathbf{q},\mathbf{u})\end{bmatrix},\quad\mathbf{y}=\mathbf{h}(\mathbf{z})=\begin{bmatrix}\mathbf{d}(\mathbf{z},\mathbf{q})\\ \mathbf{a}(\mathbf{q})\end{bmatrix}$, (75)
@@
The submatrix of $\boldsymbol{\mathcal{O}}^{(r)}$ consisting the first $\dim\mathcal{Z}$ columns has no more than $s(r+1)$ nonzero rows, so the row span of $\boldsymbol{\mathcal{O}}^{(r)}$ will always be deficient if $s(r+1)<\dim\mathcal{Z}$.
@@
Our quadrotor CLS system admits the decomposition in Proposition C.1, with:
@@
    $\mathcal{Z}$ is the space for the $6$-dimensional translational state including position and velocity $\mathbf{z}=(\mathbf{r},\mathbf{v})$;
@@
    $\mathcal{Q}$ is the space for the rotational state $\mathbf{q}$;
  3. 3.
    $\mathbf{d}$ is the $1$-dimensional range-only measurement $(s=1)$;
  4. 4.
    $\mathbf{a}(\mathbf{q})=\mathbf{q}$ is the direct measurement of the rotational state;
So applying Proposition C.1 gives $1\times(r_{*}+1)\geq\dim\mathcal{Z}=6$, as required. ∎
@@ Acknowledgements
We would like to acknowledge the sponsorship by the Natural Sciences and Engineering Research Council of Canada (NSERC) under the Discovery Grant RGPIN-2023-05148. We also acknowledge our summer student, Tommy Zhang, for contributing to flight testing and hardware interfacing of quadrotor platforms.

### v3

@@ Appendix A Derivation of closed form posterior uncertainty
whenever $\mathcal{K}^{-1}h\in L^{2}([0,T])$. Comparing Eqs. (54) and (55) shows that, for any integer $s>\alpha+1$, the condition $h\in\norm{h}_{H^{s}([0,T])}$ is sufficient for the right-hand side of Eq. (55) to converge, and in such a case an inequality of the form Eq. (52) holds.
Appendix B Derivation of the asymptotic size of the minimum eigenvalue of the STLOG
This appendix provides the derivation of Proposition 5.2. This asymptotic result relates the STLOG in Eq. (34) to the local observability index $r_{*}$ (14) defined as a bridge to rank-testing observability in Sec. 3. We consider a fixed operating point of the system $(\mathbf{x},\mathbf{u})$, and adopt several inessential simplifications. Specifically, we set $\boldsymbol{\Sigma}=\mathbf{1}$, and simply denote the order-$r$ STLOG from Eq. (34) by $\mathbf{W}^{(r)}$ without the subscript $\boldsymbol{\Sigma}^{-1}$.
We first state the main theorem of this appendix giving bounds on the minimum eigenvalue of the STLOG, then we give our proposition that bounds on this minimum eigenvalue influence local observability. Finally, we derive the relevant bounds.
@@ Appendix B Derivation of the asymptotic size of the minimum eigenvalue of the STLOG
We prove Part 1 in Proposition B.3, and derive $C_{1}(\mathbf{x},\mathbf{u}),C_{2}(\mathbf{x},\mathbf{u})$ in Propositions B.4 and B.7 respectively. ∎
While Theorem B.1 gave bounds on $\lambda_{\min}(\mathbf{W}^{(r)})$, we separately prove how such bounds can influence local weak observability of a system. Recall in Sec. 3 we stated that the system is locally weakly observable at order $r$ if $\text{rank }\boldsymbol{\mathcal{O}}^{(r)}=\dim\mathcal{X}$. We introduce the following subspace:
@@
Since $\mathcal{H}_{k-1}\supset\mathcal{H}_{k}$, this decreasing sequence of subspaces reaches $\{0\}$ precisely at the local observability index $r_{*}$. This definition leads to the following Lemma:
@@
Before incorporating the STLOG into the derivation, we rewrite it in terms of $\boldsymbol{\mathcal{O}}^{(r)}$. For any $\mathbf{v},\tilde{\mathbf{v}}\in\mathcal{X}$: [^11] In Eq. (58), if instead of the standard inner product on $\mathcal{Y}$ we use the inner product induced by $\boldsymbol{\Sigma}^{-1}$, we recover the expressions in Eqs. (27) and (34).
@@
making $\mathbf{W}^{(r)}$ positive semi-definite by construction.
@@
We first derive the upper bound in Eq. (B.1). Let $r_{*}$ be the local observability index. We only consider $\mathbf{W}^{(r)}$ with $r\geq r_{*}$ in the following, and assume $r_{*}>0$ without loss of generality.
We use the identity on analytical minima of a quadratic form over a unit sphere $\lambda_{\min}(\mathbf{W}^{(r)})=\min_{\mathbf{v}\in\mathcal{X},\ \norm{\mathbf{v}}=1}\mathbf{v}^{\top}\mathbf{W}^{(r)}\mathbf{v}$. We thus seek an appropriate unit vector $\mathbf{v}$ that demonstrates the upper bound in Eq. (B.1). Any unit vector $\mathbf{v}\in\mathcal{H}_{r_{*}-1}$ satisfies
@@
$\norm{(DL_{\mathbf{f}}^{r_{*}}\mathbf{h})\mathbf{v}}=\sigma_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}|_{\mathcal{H}_{r_{*}-1}}\right)>0,$ (61)
where $|_{\mathcal{H}_{r_{*}-1}}$ denotes the restriction to the subspace $\mathcal{H}_{r_{*}-1}$. Now we evaluate $\mathbf{v}^{\top}\mathbf{W}^{(r)}\mathbf{v}$ using our selected $\mathbf{v}$:
@@
$\lambda_{\min}(\mathbf{W}^{(r)})\leq\frac{(1+\epsilon)\Delta T^{2r_{*}+1}\sigma^{2}_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}|_{\mathcal{H}_{r_{*}-1}}\right)}{{(r_{*}!)}^{2}(2r_{*}+1)},$ (65)
@@
Next, we derive the lower bound in Eq. (56). We rewrite the STLOG into block matrix form
@@
$\alpha\sigma^{2}_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}\right)\leq\frac{\lambda_{\min}(\mathbf{W}^{(r)})}{\Delta T^{2r_{*}+1}}\leq\beta\sigma^{2}_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}|_{\mathcal{H}_{r_{*}-1}}\right),$ (72)
where $\alpha,\beta$ are constants. Thus, we proved the inequality (56) in Theorem B.1 by explicitly constructing $C_{1}(\mathbf{x},\mathbf{u}),C_{2}(\mathbf{x},\mathbf{u})$ in terms of $\boldsymbol{\mathcal{O}}^{(r_{*})}$ only. The last part of Theorem B.1 follows from the Bolzano–Weierstrass theorem.
Appendix C Derivation of the choice of STLOG order for quadrotor CLS
This appendix provides the derivation that justifies the choice of STLOG order $r=5$ for the quadrotor CLS.
@@ Appendix C Derivation of the choice of STLOG order for quadrotor CLS
$\begin{bmatrix}\dot{\mathbf{z}}\\ \dot{\mathbf{q}}\end{bmatrix}$ $=\begin{bmatrix}\mathbf{f}_{z}(\mathbf{z},\mathbf{q},\mathbf{u})\\ \mathbf{f}_{q}(\mathbf{q},\mathbf{u})\end{bmatrix},\quad\mathbf{y}=\mathbf{h}(\mathbf{z},\mathbf{q})=\begin{bmatrix}\mathbf{d}(\mathbf{z},\mathbf{q})\\ \mathbf{a}(\mathbf{q})\end{bmatrix}$, (73)
@@
The submatrix of $\boldsymbol{\mathcal{O}}^{(r)}$ consisting of the first $\dim\mathcal{Z}$ columns has no more than $s(r+1)$ nonzero rows, so the row span of $\boldsymbol{\mathcal{O}}^{(r)}$ will always be deficient if $s(r+1)<\dim\mathcal{Z}$.
@@
Applying Proposition C.1 to our quadrotor CLS gives $1\times(r_{*}+1)\geq\dim\mathcal{Z}=6$ as required, since:
@@
    $\mathcal{Z}$ is the space for the translational state, including position and velocity $\mathbf{z}=(\mathbf{r},\mathbf{v})$, giving $\dim(\mathcal{Z})=6$, and
@@
    $\mathbf{d}$ is the $1$-dimensional range-only measurement, giving $s=1$.
by the nature of how position and velocity are partially observed through range-only measurements. ∎
@@ Acknowledgements
We would like to acknowledge the sponsorship by the Natural Sciences and Engineering Research Council of Canada (NSERC) under the Discovery Grant RGPIN-2023-05148.